Could you pls check my answer?
Solve for z: |6z+30|= 48
My solution: z = 3 or z = -13
Thanks.

Answer by ewatrrr(24785)

 
Hi,
|6z+30|= 48
6z+30 = 48 z = 3
6z+30= -48 z = -13
Your solution: z = 3 or z = -13, Yes! Good work.